# Guarded Query Routing for Large Language Models

### v1

@@ Abstract
Query routing, the task to route user queries to different large language model (LLM) endpoints, can be considered as a text classification problem. However, out-of-distribution queries must be handled properly, as those could be questions about unrelated domains, queries in other languages, or even contain unsafe text. Here, we thus study a guarded query routing problem, for which we first introduce the Guarded Query Routing Benchmark (GQR-Bench), which covers three exemplary target domains (law, finance, and healthcare), and seven datasets to test robustness against out-of-distribution queries. We then use GQR-Bench to contrast the effectiveness and efficiency of LLM-based routing mechanisms (GPT-4o-mini, Llama-3.2-3B, and Llama-3.1-8B), standard LLM-based guardrail approaches (LlamaGuard and NVIDIA NeMo Guardrails), continuous bag-of-words classifiers (WideMLP, fastText), and traditional machine learning models (SVM, XGBoost). Our results show that WideMLP, enhanced with out-of-domain detection capabilities, yields the best trade-off between accuracy (88%) and speed (<4ms). The embedding-based fastText excels at speed (<1ms) with acceptable accuracy (80%), whereas LLMs yield the highest accuracy (91%) but are comparatively slow (62ms for local Llama-3.1:8B and 669ms for remote GPT-4o-mini calls). Our findings challenge the automatic reliance on LLMs for (guarded) query routing and provide concrete recommendations for practical applications. GQR-Bench will be released as a Python package — gqr. Source code: https://github.com/williambrach/gqr.
@@ 1 Introduction
Large Language Models (LLMs) are responding to millions of queries each day across a wide range of applications. To save computational resources, it is of high interest to route each query to LLMs of appropriate sizes, inference compute budgets, and domain specialization. We view this query routing problem as an instance of text classification, where the task is to classify a query into a set of suitable endpoints (e.g., LLM agents, rule-based chatbots, human chat partners). See Figure 1. While such a routing mechanism can be prepared for in-distribution queries (e.g., known languages, a pre-defined set of target domains, generally safe inputs), it must also be able to handle out-of-distribution queries (out of domain, unsupported languages, unsafe inputs). We denote this task guarded query routing.
@@
Guarded query routing is an extremely challenging task because the models cannot be prepared for all possible out-of-distribution queries. This implies that we must assume that there is no training data available for out-of-distribution queries (otherwise, they would not be out of distribution). Presumably for this reason, practitioners currently rely on LLMs to carry out the routing, such as NVIDIA NeMo Guardrails [37] or LlamaGuard [20]. However, LLM calls come with extensive computational costs, and we hypothesize that efficient text classifiers combined with an out-of-distribution detection mechanism would perform similarly well with less compute.
To test our hypothesis, we first set up a new evaluation benchmark, called Guarded Query Routing Benchmark (GQR-Bench), that simulates a guarded query routing problem. GQR-Bench is composed of three datasets that we consider our target domains (law, finance, health), as well as seven datasets from which we draw different types of out-of-distribution queries: out-of-domain queries, queries in different languages, and queries with unsafe inputs. We further provide a suggested evaluation methodology that balances in-domain accuracy and robustness against out-of-distribution queries.
Based on recent surveys in text classification [15], we then evaluate strong text classification methods such as BERT [9], fastText [23], and WideMLP [13] on GQR-Bench, testing their suitability as guarded query routers. We also test traditional text classification methods (e.g., SVM, XGBoost) on top of contextualized embeddings from a pre-trained language models. We test these methods against common strategies for standard query routing (e.g., routing via a strong LLM such as GPT-4o-mini) as well as standard guardrails to prevent unsafe inputs (e.g., NVIDIA NeMO Guardrails), which we consider as our baselines.
Our results show that efficient text classifiers are almost on par with the much more costly yet common approach of using an LLM as-a-router with WideMLP attaining 95% relative performance – while requiring orders of magnitude less time. Generally, WideMLP displays a promising trade-off between speed and accuracy, while fastText is even faster but with notable decrease in accuracy. In addition, we show that standard guardrail techniques and previous approaches to query routing do not perform well on GQR-Bench.
Generally, we find that established machine learning techniques can achieve comparable accuracy to LLMs in guarded query routing with dramatically less latency, in some cases to under 1 ms, and eliminating the costs associated with LLM calls. As such, this research challenges the automatic reliance on LLMs for query routing and provides practical alternatives for optimizing the performance-cost-latency triangle in domain-specific applications.
In summary, the main contributions of this paper are threefold:
@@
    We present a new benchmark, GQR-Bench, along with an evaluation methodology, for guarded query routing, where the routing mechanisms need to determine the correct domain, while being robust against out-of-distribution queries.
@@ 2 Related Work
The rationale behind query routing lies in its ability to improve cost efficiency while maintaining the integrity of responses [40]. This allows users to optimize the allocation of computational resources and operational expenses by directing less complex queries to smaller, more economical models, reserving more sophisticated models for tasks requiring advanced capabilities or rejecting queries identified as malicious or outside the domain of interest [18, 19, 43]. A primary application domain involves the systematic routing between models with different sizes and functional capabilities [33]. This dynamic allocation methodology enables systems to delegate less cognitively demanding tasks to models with fewer parameters, thereby preserving the computational resources of larger architectures for queries requiring enhanced precision or advanced reasoning capabilities. Studies using real-world examples show that this query routing can reduce the use of large models by up to 40% without affecting their quality [48, 10]. The theoretical foundations of query routing extend beyond mere model selection. It takes inspiration from conceptually adjacent frameworks, such as Mixture-of-Experts architectures [25, 42], which incorporate internal routing mechanisms, and Retrieval-Augmented Generation (RAG) systems. RAG systems employ contextual classification [^1] https://weaviate.io/developers/contributor-guide/contextionary/classification-benchmarks methods to select appropriate document repositories and determine whether retrieval augmentation is necessary for specific queries [5]. Despite the advancements in this field, there are still many research challenges to overcome, such as enhancing router generalization across diverse model pairings and data distributions [26, 32]; developing task-aware routing mechanisms with contextual sensitivity [47, 10]; and establishing evaluation benchmarks designed specifically for testing routing performance. In summary, while query routing is a well-studied problem in different contexts, the crucial aspect of robustness to out-of-distribution queries has not been studied so far.
@@ 4.1 Dataset Composition
As sources for out-of-distribution queries, we selected datasets that focus on toxicity (Jigsaw [22]), offensive language (OLID [51]), and hate speech (HateXplain [30]). For evaluation of capabilities to not answer queries in different languages, we employ datasets dk_hate [45] that contain offensive questions in Danish and the hate speech Slovak dataset [^2] https://huggingface.co/datasets/TUKE-KEMT/hate_speech_slovak. To evaluate no offensive/hate domains, we also employ two Q&A datasets - Machine Learning [^3] https://huggingface.co/datasets/mjphayes/machine˙learning˙questions and Web Questions [2]. It is crucial that these out-of-distribution datasets are not available for training. Instead, the query routers only see queries from these datasets at test time.
@@ 4.2 Evaluation methodology
Lastly, we consider latency and disk size efficiency metrics crucial for practical applications. Integrating high-latency models could impact inference response times and consequently degrade user experience.
We further report disk size as a robust metric comparable across different models and technical setups.
@@ 5.1 Baselines: LLMs and Embedding Similarity
Large Language Models as Query Routers: We implemented guarded query routing mechanisms using various LLMs as pre-processing filters that evaluate requests before forwarding them to the primary model. Our evaluation included GPT-4o-mini, Llama3.1-8B, and Llama3.2-3B. The employed prompt template is provided in Supplementary Material A.
Nvidia NeMo Guardrails: This framework [37] functions as a protective guardrail for deployed LLMs by utilizing a pre-processing evaluation mechanism to assess input conformity with established safety parameters. We evaluated the NeMo guardrail system with multiple LLM gatekeepers—specifically GPT-4o-mini, Llama3.1-8B, and Llama3.2-3B. The guardrail configuration is provided in Supplementary Material A.
@@ 5.2 Efficient Text Classifiers
WideMLP [13] has attained best performance in a comprehensive survey [15], even surpassing LLMs. WideMLP operates on bag-of-words input representations and employs a small number of wide hidden layers. In particular, we employed one hidden layer with 1024 hidden units. The initial hidden layer is implemented efficiently through an embedding table, and the pooling over the sequential dimension is weighted by the tokens’ IDF scores (as in TF-IDF).
We adapted WideMLP for the guarded query router by making two key changes: First, we trained the model with binary cross-entropy instead of categorical cross-entropy to avoid overconfident mispredictions caused by the softmax. Second, we employed a threshold on the confidence scores to detect out-of-domain queries: When the predicted probability for all classes falls below this threshold, we consider the current example out-of-distribution. We systematically evaluated multiple threshold configurations (0.5, 0.7, 0.9, and 0.99) to determine the optimal balance between domain classification accuracy and out-of-domain detection. These changes are inspired by previous work on text classification under the presence of OOD examples [44], which further investigates class-specific thresholds obtained through a risk reduction technique. However, follow-up work has found that these class-specific thresholds obtained through risk reduction are not more effective than a correctly set global threshold [14] – which is why we do not employ risk reduction here.
fastText [4] has extended word2vec [31] by a fall-back for out-of-vocabulary words through bags of character n-grams, enabling effective handling of out-of-vocabulary terms. For classification, fastText employs pre-trained word embeddings and trains a logistic regression on top [23]. The pre-trained embeddings come in a large variety of languages, each trained mainly on the Wikipedia corpus of the respective language.
To adapt fastText for guarded query routing, we employed a One-vs-Rest approach, whereby we train one binary classifier per domain that produces a binary prediction (0 or 1) to predict domain membership. When all domain-specific classifiers return 0 for a given query, the system identifies it as out-of-distribution and rejects it accordingly. Hyperparameter selection was handled using fastText’s autotune feature, with a time budget set to 300 seconds per classifier to find optimal settings based on validation performance.
@@ 6 Results
Table 2: Main Results. In-distribution (ID) accuracy, out-of-distribution (OOD) accuracy, and GQR-Score as combined score between ID and OOD accuracy. Best methods per category of approaches highlighted in bold. The highest GQR-Score is achieved by Llama3.1:8b, closely followed by GPT-4o-mini and then WideMLP.
Model ID Acc. OOD Acc. GQR-Score
Baselines: Routing based on large language models
Llama3.1:8b 95.66 88.00 91.67
GPT-4o-mini 95.70 87.85 91.61
Llama3.2:3b 26.37 99.82 41.72
Baselines: Embedding similarity approaches
all-MiniLM-L6-v2 + S.R. 90.00 42.45 57.69
bge-small-en-v1.5 + S.R. 90.70 35.39 50.91
Continuous bag-of-words classifiers
fastText 95.80 68.85 80.12
WideMLP(t=0.99) 84.49 91.25 87.74
WideMLP(t=0.90) 90.91 80.39 85.33
WideMLP(t=0.75) 93.67 73.33 82.26
Fine-tuned encoder-only language models
ModernBERT-base (t=0.99) 99.94 19.74 32.97
BERT-base (t=0.99) 99.90 19.80 33.05
Sentence embeddings + traditional classifiers
bge-small-en-v1.5 + SVM 99.42 71.15 82.94
bge-small-en-v1.5 + XGBoost 98.78 68.70 81.04
all-MiniLM-L6-v2 + SVM 86.06 63.14 72.84
all-MiniLM-L6-v2 + XGBoost 92.93 68.73 79.02
all-MiniLM-L12-v2 + MLP 95.17 59.60 73.23
TF-IDF + XGBoost 34.76 84.49 49.26
TF–IDF + SVM 34.36 55.31 42.39
@@
NeMo Guardrails + Llama3.2:8B 99.04
@@
Table 4 provides a breakdown of OOD performance across the different datasets, allowing us to assess model robustness to various types of out-of-distribution queries.
@@
Table 4: Breakdown of OOD accuracy across all considered datasets, and each model’s unweighted average accuracy. The best average accuracy is highlighted per category of model. Manually prompted Llama3.2:3b performs best, followed by WideMLP.
Model Jigsaw OLID HateXplain dkhate TUKE SK Web Q ML Q Average
Baselines: Routing based on large language models
Llama-Guard-3-8B 27.07 24.77 93.28 5.17 7.51 0.10 0.00 22.56
Llama-Guard-3-1B 51.40 61.40 91.47 12.77 20.13 2.31 0.00 34.21
Llama3.2:3b 99.69 99.88 99.98 100.00 100.00 99.16 100.00 99.82
Llama3.1:8b 94.43 93.60 97.99 95.74 97.60 90.55 46.09 88.00
NeMo Guardrails + Llama3.2:3B 61.42 59.65 43.15 61.09 67.88 1.67 0.00 42.12
NeMo Guardrails + Llama3.1:8B 51.99 36.40 20.83 10.33 27.11 0.00 0.00 20.95
NeMo Guardrails + GPT-4o-mini 98.26 94.19 99.78 91.49 96.14 57.19 79.69 88.11
GPT-4o-mini 94.71 93.49 98.10 94.53 98.02 90.80 45.31 87.85
Baselines: Embedding similarity approaches
all-MiniLM-L6-v2 + Semantic Router (s=5, t=0.5) 22.96 31.74 36.71 39.51 20.33 96.70 49.22 42.45
bge-small-en-v1.5 + Semantic Router (s=5, t=0.5) 15.15 28.95 32.67 31.91 12.41 95.42 31.25 35.39
Continuous bag-of-words classifiers
fastText 74.46 61.51 54.46 74.77 83.11 70.37 63.28 68.85
WideMLP (t=0.99) 93.83 93.49 91.00 86.93 80.60 99.16 93.75 91.25
WideMLP (t=0.90) 87.87 83.26 77.56 71.73 56.93 95.57 89.84 80.39
WideMLP (t=0.75) 84.04 76.74 70.48 57.45 47.34 92.91 84.38 73.33
Fine-tuned encoder-only language models
ModernBERT-base (t=0.99) 27.10 17.91 18.06 10.33 2.50 62.30 0.00 19.74
BERT-base-multilingual-cased (t=0.99) 20.91 28.26 25.44 25.84 30.87 7.28 0.00 19.80
Sentence embeddings + traditional classifiers
bge-small-en-v1.5 + SVM 77.47 75.00 63.81 61.40 63.82 59.69 96.88 71.15
bge-small-en-v1.5 + XGBoost 81.95 68.26 72.15 47.72 59.02 58.81 92.97 68.70
all-MiniLM-L6-v2 + SVM 59.61 71.74 61.63 37.99 34.62 81.89 94.53 63.14
all-MiniLM-L6-v2 + XGBoost 47.57 77.44 53.14 57.45 60.17 95.47 89.84 68.73
all-MiniLM-L12-v2 + MLP 74.77 80.47 85.59 56.23 18.87 68.45 32.81 59.60
TF–IDF + SVM 24.58 26.16 21.72 75.38 96.98 54.87 87.50 55.31
TF–IDF + XGBoost 58.31 67.44 66.40 100.00 99.90 99.36 100.00 84.49
@@
Model Size on disk MB Batch Size
@@
Table 5 provides an overview of latency across several batch sizes. LLM-based approaches consistently exhibit the highest latencies. GPT-4o-mini shows the highest single-query latency (because of REST API call) at 0.667 seconds, while Llama models range from 0.042 to 0.083 seconds per query. Batch processing improves efficiency the LLMs. Semantic Router yields a moderate latency (0.035 seconds per query) with no improvements through batching. FastText shows exceptional efficiency with latencies of approximately 0.00007-0.00009 seconds per query, roughly three orders of magnitude faster than LLM approaches. SVM and XGBoost models paired with various embedding approaches show excellent efficiency. Specifically, the embedding-based XGBoost models benefit substantially from batch processing, improving from 0.022 seconds per query to as low as 0.00035 seconds per query. All experiments were conducted on a single server with specifications: 2× ASUS GeForce RTX 4090 GPUs, AMD Threadripper PRO 7965WX (24-Core) processor, and 256GB RAM.
Figure 2: Efficiency-effectiveness tradeoff showing harmonic mean performance versus latency (log scale). Models in the upper-left quadrant offer optimal balance between high performance and low latency.
@@
Model size on disk varies substantially across routing models (see Table 5), with important implications for deployment scenarios where storage and memory are constrained (e.g., edge devices). LLM-based approaches require the most storage, with Llama models ranging from approximately 2 GB to 4.9 GB. In contrast, traditional classification approaches based on TF-IDF require only between 4 and 11 MB. Models based on all-MiniLM-L12-v2 require about 100 MB, and the quantized MLP uses about 21.8 MB. Our one-vs-rest fastText approach uses 221 MB of disk size.

### v3

@@ Abstract
Query routing, the task to route user queries to different large language model (LLM) endpoints, can be considered as a text classification problem. However, out-of-distribution queries must be handled properly, as those could be about unrelated domains, queries in other languages, or even contain unsafe text. Here, we thus study a guarded query routing problem, for which we first introduce the Guarded Query Routing Benchmark (GQR-Bench, released as Python package gqr), covers three exemplary target domains (law, finance, and healthcare), and seven datasets to test robustness against out-of-distribution queries. We then use GQR-Bench to contrast the effectiveness and efficiency of LLM-based routing mechanisms (GPT-4o-mini, Llama-3.2-3B, and Llama-3.1-8B), standard LLM-based guardrail approaches (LlamaGuard and NVIDIA NeMo Guardrails), continuous bag-of-words classifiers (WideMLP, fastText), and traditional machine learning models (SVM, XGBoost). Our results show that WideMLP, enhanced with out-of-domain detection capabilities, yields the best trade-off between accuracy (88%) and speed (<4ms). The embedding-based fastText excels at speed (<1ms) with acceptable accuracy (80%), whereas LLMs yield the highest accuracy (91%) but are comparatively slow (62ms for local Llama-3.1:8B and 669ms for remote GPT-4o-mini calls). Our findings challenge the automatic reliance on LLMs for (guarded) query routing and provide concrete recommendations for practical applications. Source code is available: https://github.com/williambrach/gqr.
@@ 1 Introduction
Large Language Models (LLMs) are responding to millions of queries each day. To save computational resources, it is of high interest to route each query to LLMs of appropriate sizes, inference compute budgets, and domain specialization. We view this query routing problem as an instance of text classification, where the task is to classify a query into a set of suitable endpoints (e.g., LLM agents, rule-based chatbots, human chat partners). See Figure 1.
While such a routing mechanism can be prepared for in-distribution queries (e.g., known languages, a pre-defined set of target domains, generally safe inputs), it must also deal with out-of-distribution queries, such as queries in other domains, languages, or even unsafe inputs.
@@
We denote this task as guarded query routing. Guarded query routing is an extremely challenging task because the models cannot be prepared for all possible out-of-distribution queries. This implies that we must assume that there is no training data available for out-of-distribution queries (otherwise, they would not be out of distribution). Presumably for this reason, practitioners currently rely on LLMs to carry out the routing, such as NVIDIA NeMo Guardrails [38] or LlamaGuard [21]. However, LLM calls come with extensive computational costs, and we hypothesize that efficient text classifiers combined with an out-of-distribution detection mechanism would perform similarly well with less compute.
To test our hypothesis, we first set up a new evaluation benchmark, called Guarded Query Routing Benchmark (GQR-Bench), that simulates a guarded query routing problem. GQR-Bench is composed of three datasets that we consider our target domains (law, finance, health), as well as seven datasets from which we draw different types of out-of-distribution queries: out-of-domain queries, queries in different languages, and queries with unsafe inputs. We further provide a suggested evaluation methodology that balances in-domain accuracy and robustness against out-of-distribution queries. Based on recent surveys in text classification [16], we then evaluate strong text classification methods such as BERT [10], fastText [24], and WideMLP [14] on GQR-Bench, testing their suitability as guarded query routers. We also test traditional text classification methods (e.g., SVM, XGBoost) on top of contextualized embeddings from a pre-trained language model. We test these methods against common strategies for standard query routing (e.g., routing via a strong LLM such as GPT-4o-mini) as well as standard guardrails to prevent unsafe inputs (e.g., NVIDIA NeMO Guardrails), which we consider as our baselines. We show that efficient text classifiers can rival costly LLM-based routers for guarded query routing. WideMLP achieves 95% of the LLM’s relative performance while being orders of magnitude faster (<1ms latency) and eliminating API costs. We also demonstrate that standard guardrails and prior routing methods fail on our GQR-Bench, underscoring the task’s difficulty. This work challenges the default reliance on LLMs for routing, offering a practical, high-performance alternative.
In summary, the main contributions of this paper are:
@@
    We present a new benchmark, GQR-Bench is released as a Python package (https://pypi.org/project/gqr/), along with an evaluation methodology, for guarded query routing, where the routing mechanisms need to determine the correct domain, while being robust against out-of-distribution queries.
@@ 2 Related Work
The rationale behind query routing lies in its ability to improve cost efficiency while maintaining the integrity of responses [41]. This allows users to optimise the allocation of computational resources and operational expenses by directing less complex queries to smaller, more economical models, reserving more sophisticated models for tasks requiring advanced capabilities or rejecting queries identified as malicious or outside the domain of interest [19, 20, 44]. A primary application domain involves the systematic routing between models with different sizes and functional capabilities, often using contrastive learning to differentiate between the strengths of various models [34, 6]. This dynamic allocation methodology enables systems to delegate less cognitively demanding tasks to models with fewer parameters, thereby conserving the computational resources of larger architectures for queries that require enhanced precision or advanced reasoning capabilities. Studies using real-world examples show that this query routing can reduce the use of large models by up to 40% without affecting their quality [50, 11]. The theoretical foundations of query routing extend beyond mere model selection. It takes inspiration from conceptually adjacent frameworks, such as Mixture-of-Experts architectures [26, 43], which incorporate internal routing mechanisms, and Retrieval-Augmented Generation (RAG) systems. RAG systems employ contextual classification [^1] https://weaviate.io/developers/contributor-guide/contextionary/classification-benchmarks methods to select appropriate document repositories and determine whether retrieval augmentation is necessary for specific queries [5]. Despite the advancements in this field, there are still many research challenges to overcome, such as enhancing router generalization across diverse model pairings and data distributions [27, 33]; developing task-aware routing mechanisms with contextual sensitivity [49, 11]; and establishing evaluation benchmarks designed specifically for testing routing performance. In summary, while query routing is a well-studied problem in different contexts, the crucial aspect of robustness to out-of-distribution queries has not been studied so far.
@@ 4.1 Dataset Composition
As sources for out-of-distribution queries, we selected datasets that focus on toxicity (Jigsaw [23]), offensive language (OLID [53]), and hate speech (HateXplain [31]). For evaluation of capabilities to not answer queries in different languages, we employ datasets dk_hate [46] that contain offensive questions in Danish and the hate speech Slovak dataset [^2] https://huggingface.co/datasets/TUKE-KEMT/hate_speech_slovak. To evaluate no offensive/hate domains, we also employ two Q&A datasets - Machine Learning [^3] https://huggingface.co/datasets/mjphayes/machine˙learning˙questions and Web Questions [2]. These out-of-distribution datasets mustn’t be available for training. Instead, the query routers only see queries from these datasets at test time.
@@ 4.2 Evaluation methodology
Lastly, we consider latency and disk size efficiency metrics crucial for practical applications. Integrating high-latency models could impact inference response times and consequently degrade user experience. We further report disk size as a robust metric comparable across different models and technical setups.
@@ 5.1 Baselines: LLMs and Embedding Similarity
Large Language Models as Query Routers: We implemented guarded query routing mechanisms using various LLMs as pre-processing filters that evaluate requests before forwarding them to the primary model. Our evaluation included GPT-4o-mini, Llama3.1-8B, and Llama3.2-3B. The employed prompt templates are provided in the arXiv version of this paper [47].
Nvidia NeMo Guardrails: This framework [38] functions as a protective guardrail for deployed LLMs by utilizing a pre-processing evaluation mechanism to assess input conformity with established safety parameters. We evaluated the NeMo guardrail system with multiple LLM gatekeepers—specifically GPT-4o-mini, Llama3.1-8B, and Llama3.2-3B. The employed guardrail configurations are provided in the arXiv version of this paper [47].
@@ 5.2 Efficient Text Classifiers
WideMLP [14] has attained best performance in a comprehensive survey [16], even surpassing LLMs. WideMLP operates on bag-of-words input representations and employs a small number of wide hidden layers. We used the default hyperparameters for WideMLP. This architecture consists of one hidden layer with 1024 hidden units, followed by a dropout layer with a rate of 0.5. The initial hidden layer is implemented efficiently through an embedding table, and the pooling over the sequential dimension is weighted by the tokens’ IDF scores (as in TF-IDF). We adapted WideMLP for the guarded query router by making two key changes: First, we trained the model with binary cross-entropy instead of categorical cross-entropy to avoid overconfident mispredictions caused by the softmax. Second, we employed a threshold on the confidence scores to detect out-of-domain queries: When the predicted probability for all classes falls below this threshold, we consider the current example out-of-distribution. We systematically evaluated multiple threshold configurations (0.5, 0.7, 0.9, and 0.99) to determine the optimal balance between domain classification accuracy and out-of-domain detection. These changes are inspired by previous work on text classification under the presence of OOD examples [45], which further investigates class-specific thresholds obtained through a risk reduction technique. However, follow-up work has found that these class-specific thresholds obtained through risk reduction are not more effective than a correctly set global threshold [15] – which is why we do not employ risk reduction here.
fastText [4] has extended word2vec [32] by a fall-back for out-of-vocabulary words through bags of character n-grams, enabling effective handling of out-of-vocabulary terms. For classification, fastText employs pre-trained word embeddings and trains a logistic regression on top [24]. The pre-trained embeddings come in a large variety of languages, each trained mainly on the Wikipedia corpus of the respective language. To adapt fastText for guarded query routing, we employed a One-vs-Rest approach, whereby we train one binary classifier per domain that produces a binary prediction (0 or 1) to predict domain membership. When all domain-specific classifiers return 0 for a given query, the system identifies it as out-of-distribution and rejects it accordingly. Hyperparameter selection was handled using fastText’s autotune feature, with a time budget set to 300 seconds per classifier to find optimal settings based on validation performance.
@@ 5.3 Sentence Embeddings with Traditional Classifiers
Table 2: This table shows the performance of various models on in-distribution (ID) and out-of-distribution (OOD) datasets, which include both unsafe and out-of-domain data. The table presents OOD and ID accuracies for each dataset, along with three summary metrics: Unsafe Avg (the average accuracy on the five unsafe content datasets), OOD Accuracy, and GQR Score. The best score in each column is highlighted. While the manually prompted Llama3.2:3B model achieves the highest OOD accuracy, the overall best model according to the GQR score is the manually prompted Llama3.1:8B, followed by the manually prompted GPT-4o-mini and the WideMLP. Standard guardrail methods were not applicable to classify in-domain queries, which is marked by ‘—’.
Model Jigsaw OLID HateXplain dkhate TUKE SK Web Q ML Q Unsafe Avg. ID Acc. OOD Acc. GQR score
Baselines: Standard guardrail methods
Llama-Guard-3-1B 51.40 61.40 91.47 12.77 20.13 2.31 0.00 47.43 34.21
Llama-Guard-3-8B 27.07 24.77 93.28 5.17 7.51 0.10 0.00 31.56 22.56
NeMo Guardrails + Llama3.2:3B 61.42 59.65 43.15 61.09 67.88 1.67 0.00 58.64 58.64
NeMo Guardrails + Llama3.1:8B 51.99 36.40 20.83 10.33 27.11 0.00 0.00 29.33 29.33
NeMo Guardrails + GPT-4o-mini 98.26 94.19 99.78 91.49 96.14 57.19 79.69 95.97 95.97
Baselines: Embedding similarity approaches
all-MiniLM-L6-v2 + Semantic Router (s=5, t=0.5) 22.96 31.74 36.71 39.51 20.33 96.70 30.25 49.22 90.00 42.45 57.69
bge-small-en-v1.5 + Semantic Router (s=5, t=0.5) 15.15 28.95 32.67 31.91 12.41 95.42 31.25 24.22 90.70 35.39 50.91
Baselines: Routing based on large language models
Llama3.2:3B 99.69 99.88 99.98 100.00 100.00 99.16 100.00 99.91 26.37 99.82 41.72
Llama3.1:8B 94.43 93.60 97.99 95.74 97.60 90.55 46.09 95.87 95.66 88.00 91.67
GPT-4o-mini 94.71 93.49 98.10 94.53 98.02 90.80 45.31 95.77 95.70 87.85 91.61
Continuous bag-of-words classifiers
fastText 74.46 61.51 54.46 74.77 83.11 70.37 63.28 69.66 95.80 68.85 80.12
WideMLP (t=0.99) 93.83 93.49 91.00 86.93 80.60 99.16 93.75 89.17 84.49 91.25 87.74
WideMLP (t=0.90) 87.87 83.26 77.56 71.73 56.93 95.57 89.84 75.47 90.91 80.39 85.33
WideMLP (t=0.75) 84.04 76.74 70.48 57.45 47.34 92.91 84.38 67.21 93.67 73.33 82.26
Fine-tuned encoder-only language models
ModernBERT-base (t=0.99) 27.10 17.91 18.06 10.33 2.50 62.30 0.00 15.18 99.94 19.74 32.97
BERT-base-multilingual-cased (t=0.99) 20.91 28.26 25.44 25.84 30.87 7.28 0.00 26.26 99.90 19.80 33.05
Sentence embeddings + traditional classifiers
bge-small-en-v1.5 + SVM 77.47 75.00 63.81 61.40 63.82 59.69 96.88 68.30 99.42 71.15 82.94
bge-small-en-v1.5 + XGBoost 81.95 68.26 72.15 47.72 59.02 58.81 92.97 65.82 98.78 68.70 81.04
all-MiniLM-L6-v2 + SVM 59.61 71.74 61.63 37.99 34.62 81.89 94.53 53.12 86.06 63.14 72.84
all-MiniLM-L6-v2 + XGBoost 47.57 77.44 53.14 57.45 60.17 95.47 89.84 59.15 92.93 68.73 79.02
all-MiniLM-L12-v2 + MLP 74.77 80.47 85.59 56.23 18.87 68.45 32.81 63.19 95.17 59.60 73.23
TF–IDF + SVM 24.58 26.16 21.72 75.38 96.98 54.87 87.50 48.96 37.76 55.31 49.26
TF–IDF + XGBoost 58.31 67.44 66.40 100.00 99.90 99.36 100.00 78.41 34.76 84.49 42.39
@@ 6 Results
NeMo Guardrails + Llama3.1:8B 99.04
@@
Table 2 summarises model robustness to out-of-distribution (OOD) queries. We report average performance across unsafe and out-of-domain datasets.
@@
Model Size on disk Batch Size
@@
Table 4 provides an overview of latency across several batch sizes. LLM-based approaches consistently exhibit the highest latencies. GPT-4o-mini shows the highest single-query latency (because of REST API call) at 0.667 seconds, while Llama models range from 0.042 to 0.083 seconds per query. Batch processing improves efficiency of the LLMs. Semantic Router yields a moderate latency (0.035 seconds per query) with no improvements through batching. FastText shows exceptional efficiency with latencies of approximately 0.00007-0.00009 seconds per query, roughly three orders of magnitude faster than LLM approaches. SVM and XGBoost models paired with various embedding approaches show excellent efficiency. Specifically, the embedding-based XGBoost models benefit substantially from batch processing, improving from 0.022 seconds per query to as low as 0.00035 seconds per query. All experiments were conducted on a single server with specifications: 2× ASUS GeForce RTX 4090 GPUs, AMD Threadripper PRO 7965WX (24-Core) processor, and 256GB RAM.
Figure 2: Efficiency-effectiveness tradeoff showing harmonic mean performance versus latency in seconds (log scale). Models in the upper-left quadrant offer an optimal balance between high performance and low latency.
@@
Model size on disk varies substantially across routing models (see Table 4), with important implications for deployment scenarios where storage and memory are constrained (e.g., edge devices). LLM-based approaches require the most storage, with Llama models ranging from approximately 2 GB to 4.9 GB. In contrast, traditional classification approaches based on TF-IDF require only between 4 and 11 MB. Models based on all-MiniLM-L12-v2 require about 100 MB, and the quantized MLP uses about 21.8 MB. Our one-vs-rest fastText approach uses 221 MB of disk space.